An iron ring has a X-section of 2 cm2 and a mean diameter of 20 cm. An air-gap of 0.4 mm has been cut across the section of the ring. The ring is wound with a coil of 300 turns . If the total magnetic flux is 0.20 mWb. The relative permeability of iron is 1000. Determine the current flowing through the coil


1
Expert's answer
2021-02-17T14:43:24-0500

Answer

Given data

Φ = 0.24 mWb;

A = 2cm2

= 0.0003 m2 ;

 B = Φ/A

= "1Wb\/m^2"

AT for iron ring = H × l

="(B\/\u00b50 \u00b5r ) \u00d7 l"

= "(\frac{1}{12.57\times10^{−7 }× 1000} )\times0.20\\=1592"

AT for air-gap

= H × l

= (B/µ0) × l

"(\frac{1}{12.57× 10^{−7}}) × 0.4 × 10^{−3}\\ = 318"

total AT provided = 300 × I = 300 I

"300I=1592+318"

"I=6.37A"